### v2

# Hierarchical Position Embedding of Graphs with Landmarks and Clustering for Link Prediction

Abstract.
Learning positional information of nodes in a graph is important for link prediction tasks. We propose a representation of positional information using representative nodes called landmarks. A small number of nodes with high degree centrality are selected as landmarks which serve as reference points for the nodes’ positions. We justify this selection strategy for well-known random graph models and derive closed-form bounds on the average path lengths involving landmarks. In a model for power-law graphs, we prove that landmarks provide asymptotically exact information on inter-node distances. We apply theoretical insights to real-world graphs and propose Hierarchical Position embedding with Landmarks and Clustering (HPLC). HPLC combines landmark selection and graph clustering, i.e., the graph is partitioned into densely connected clusters in which nodes with the highest degree are selected as landmarks. HPLC leverages the positional information of nodes based on landmarks at various levels of hierarchy such as nodes’ distances to landmarks, inter-landmark distances and hierarchical grouping of clusters. Experiments show that HPLC achieves state-of-the-art performances of link prediction on various datasets in terms of HIT@K, MRR, and AUC. The code is available at https://github.com/kmswin1/HPLC.
Link Prediction, Network Science, Graph Neural Networks
copyright: acmcopyrightjournalyear: 2024copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singapore.booktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporeprice: 15.00isbn: 979-8-4007-0171-9/24/05doi: 10.1145/3589334.3645372conference: The Web Conference; May 13–17, 2024; SINGAPOREprice: 15.00isbn: 979-8-4007-0171-9/24/05ccs: Computing methodologies Learning latent representations
1. Introduction
Graph Neural Networks are foundational tools for various graph related tasks such as node classification (Gilmer et al., 2017; Kipf and Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018), link prediction (Adamic and Adar, 2003; Kipf and Welling, 2016; Zhang and Chen, 2018), graph classification (Xu et al., 2019), and graph clustering (Parés et al., 2017). In this paper, we focus on the task of link prediction using GNNs.
Message passing GNNs (MPGNNs) (Gilmer et al., 2017; Kipf and Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018) have been successful in learning structural node representations through neighborhood aggregation. However, standard MPGNNs achieved subpar performances on link prediction tasks due to their inability to distinguish isomorphic nodes. In order to tackle isomorphism, subgraph-based methods (Zhang and Chen, 2018; Pan et al., 2022) were proposed for link prediction based on dynamic node embeddings from enclosing subgraphs. Labeling methods (Li et al., 2020; Zhang et al., 2021) assign node labels to identify each node via hashing or aggregation functions. An important line of research is the position-based methods  (You et al., 2019; Dwivedi et al., 2020; Kreuzer et al., 2021; Wang et al., 2022) which propose to learn and utilize the positional information of nodes.
The positional information of nodes is useful for link prediction because, e.g., isomorphic nodes can be distinguished by their positions, and the link formation between a node pair may be related to their relative distance (Srinivasan and Ribeiro, 2020). The position of a node can be defined as its distances relative to other nodes. A relevant result by Bourgain (Bourgain, 1985) states that inter-node distances can be embedded into Euclidean space with small errors. Linial (Linial et al., 1995) and P-GNN (You et al., 2019) focus on constructing Bourgain’s embeddings based on the nodes’ distances to random node sets. Another line of positional encoding used the eigenvectors of graph Laplacian (Dwivedi and Bresson, 2021; Kreuzer et al., 2021) based on spectral graph theory. However, the aforementioned methods have convergence issues and may not scale well for large or dense graphs (Wang et al., 2022). For example, Laplacian methods have stability issues (Wang et al., 2022) and may not outperform methods based on structural features (Zhang and Chen, 2018). Thus, there is significant room for improving both performance and scalability of positional node embedding for link prediction.
In this paper, we propose an effective and efficient representation of the positional information of nodes. We select a small number of representative nodes called landmarks and impose hierarchy on the graph by associating each node with a landmark. Each node computes distances to the landmarks, and each landmark computes distances to other landmarks. Such distance information is combined so as to represent nodes’ positional information. Importantly, we select landmarks and organize the graph in a principled way, unlike previous methods using random selection, e.g., (Bourgain, 1985; Linial et al., 1995; You et al., 2019).
The key question is how to select landmarks. We advocate the selection based on degree centrality which is motivated from the theory of network science. In network models with preferential attachment (PA), nodes with high degrees called hubs are likely to emerge (Barabási and Albert, 1999). PA is a process such that, if a new node joins the graph, it is more likely to connect to nodes with higher degrees. PA induces the power-law distribution of node degrees, and the network exhibits scale-free property (Barabási and Albert, 1999). Hubs are abundant in social/citation networks and World Wide Web, and play a central role in characterizing the network.
We provide a theoretical justification of landmark selection based on degree centrality for a well-known class of random graphs (Erdős et al., 1960; Barabási and Albert, 1999). We show that the inter-node distances are well-represented by the detour via landmarks. In networks with preferential attachment, we show that the strategy of choosing high-degree nodes as landmarks is asymptotically optimal in the following sense: the minimum distance among the detours via landmarks is asymptotically equal to the shortest-path distance. We show even a small number of landmarks relative to network size, e.g., $O(\log N)$, suffice to achieve optimality. This proves that the hub-type landmarks offer short paths to nodes, manifesting small-world phenomenon (Watts and Strogatz, 1998; Barabási and Albert, 1999). In addition, we show that in models where hubs are absent, one can reduce the detour distance by selecting a higher number of landmarks.
Motivated by the theory, we propose Hierarchical Position embedding with Landmarks and Clustering (HPLC). HPLC partitions a graph into $O(\log N)$ clusters which are locally dense, and appoints the node with the highest degree in the cluster as the landmark. Our intention is to bridge gap between theory and practice: hubs may not be present in real-world graphs. Thus, it is important to distribute the landmarks evenly over the graph through clustering, so that nodes can access nearby/local landmarks. Next, we form a graph of higher hierarchy, i.e., the graph of landmarks, and compute its Laplacian encoding based on the inter-landmark distances. The encoding is assigned as a membership to the nodes belonging to the cluster so as to learn positional features at the cluster level. We further optimize our model using the encoding based on hierarchical grouping of clusters. The computation of HPLC can be mainly done during preprocessing, incurring low computational costs. Our experiments on 7 datasets with 16 baseline methods show that HPLC achieves state-of-the-art performances, demonstrating its effectiveness especially over prior position-/distance-based methods.
Our contributions are summarized as follows: 1) we propose HPLC, a highly efficient algorithm for link prediction using hierarchical position embedding based on landmarks combined with graph clustering; 2) building upon network science, we derive closed-form bounds on average path lengths via detours for well-known random graphs which, to our belief, are important theoretical findings; 3) we conduct extensive experiments to show that HPLC achieves state-of-the-art performances in most cases.
2. Random Graphs with Landmarks
2.1. Notation
We consider undirected graph $G=(V,E)$ where $V$ and $E\subseteq V\times V$ denote the set of vertices and edges, respectively. Let $N$ denote the number of nodes, or $N=|V|$. $\mathbf{A}\in\mathbb{R}^{N\times N}$ denotes the adjacency matrix of $G$. $d(v,u)$ denotes the geodesic (shortest-path) distance between $v$ and $u$. Node attributes are defined as $X=\{x_{1},...,x_{N}\}$ where $x_{i}\in\mathbb{R}^{n}$ denotes the feature vector of node $i$. We consider methods of embedding $X$ into the latent space as vectors $Z=\{z_{1},...,z_{N}\},z_{i}\in\mathbb{R}^{m}$. We study the node-pair-level task of estimating the link probability between $u$ and $v$ from $z_{u}$ and $z_{v}$.
2.2. Representation of Positions using Landmarks
The distances between nodes provide rich information on the graph structures. For example, a connected and undirected graph can be represented by a finite metric space with its vertex set and the inter-node distances, and the position of a node can be defined as its relative distances to other nodes. However, computing and storing shortest paths for all pairs of nodes incur high complexity.
Instead, we select a small number of representative nodes called landmarks. The landmarks are denoted by $\lambda_{1},\cdots,\lambda_{K}\in V$ where $K$ denotes the number of landmarks. For node $v$, we define $K$-tuple of distances to landmarks as distance vector (DV):
$D(v):=\left(d(v,\lambda_{1}),d(v,\lambda_{2}),\cdots,d(v,\lambda_{K})\right)$
$D(v)$ can be used to represent the (approximate) position of node $v$. If the positions of $u$ and $v$ are to be used for link prediction, one can combine $D(u)$ and $D(v)$ (or the embeddings thereof) to estimate the probability of link formation.
The key question is, how much information $D(\cdot)$ has on the inter-node distances within the graph, so that positional information can be (approximately) recovered from $D(\cdot)$. For example, from triangle inequality, the distance between nodes $u$ and $v$ are bounded as
$d(u,v)\leq\min_{i=1,\ldots,K}[d(u,\lambda_{i})+d(\lambda_{i},v)]$
which states that the detour via landmarks, i.e., from $u$ to $\lambda_{i}$ to $v$, is longer than $d(u,v)$, but the shortest detour, or the minimum component of $D(u)+D(v)$, may provide a good estimate of $d(u,v)$. Indeed, it was empirically shown that the shortest detour via base nodes (landmarks) can be an excellent estimate of distances in Internet (Ng and Zhang, 2002).
We examine the achievable accuracy of distance estimates based on the detours via landmarks. Previous works on distance estimates using auxiliary nodes relied on heuristics (Potamias et al., 2009) or complex algorithms for arbitrary graphs (Linial et al., 1995; You et al., 2019). However, real-world graphs have structure, e.g., preferential attachment or certain distribution of node degrees. Presumably, one can exploit such structure to devise efficient algorithms. The key design questions are: how to select good landmarks, and how many of them? Drawing upon the theory of network science, we analyze a well-known class of random graphs, derive the detour distances via landmarks, and glean design insights from the analysis.
(1) $P(L_{ij}>s)=\exp\bigg{[}-\frac{h_{i}h_{j}}{\beta N\langle h^{2}\rangle}\cdot\langle h^{2}\rangle_{Q}\cdot K(N)\cdot(s-1)\bigg{(}\frac{\langle h^{2}\rangle N}{\beta}\bigg{)}^{s-1}\bigg{]},\quad s=1,2,\cdots$
(2) $\bar{l}\leq\dfrac{-2\langle\log h\rangle-\log\left(\langle h^{2}\rangle_{Q}\cdot K(N)\right)+\log(N\beta\langle h^{2}\rangle)+\log\log\left(\frac{N\langle h^{2}\rangle}{\beta}\right)-\gamma}{\log N+\log\langle h^{2}\rangle-\log\beta}+\frac{1}{2}$
2.3. Path lengths via Landmarks in random graphs
The framework by (Fronczak et al., 2004) provides a useful tool for analyzing path lengths for a wide range of classes of random graphs as follows. Consider a random graph where the probability of the existence of an edge for node $i$ and $j$ denoted by $q_{ij}$ is given by
(3) $q_{ij}=\frac{h_{i}h_{j}}{\beta}$
where $h_{i}$ is tag information of node $i$, e.g., the connectivity or degree of the node, and $\beta$ is a parameter depending on the network model. From the continuum approximation (Albert and Barabási, 2002) for large graphs, tag information $h$ is regarded as a continuous random variable (RV) with distribution $\rho(\cdot)$. For some measurable function $f$, let $\langle f(h)\rangle$ denote the expectation of $f(h)$ of a node chosen at random, and $\langle f(h)\rangle_{Q}$ denote the expectation of $f(h)$ of landmarks chosen under some distribution $Q$.
Theorem 1.
Let $L_{ij}$ denote the random variable representing the minimum path length from node $i$ to $j$ among the detours via $K(N)$ landmarks. The landmarks are chosen i.i.d. according to distribution $Q$. Asymptotically in $N$, $P(L_{ij}>s)$ is given by (1).
The proof of Theorem 1 is in Appendix A.1. In (1), the design parameters are $\langle h^{2}\rangle_{Q}$ and $K(N)$ which are related to what kind of landmarks are chosen, and how many of them, respectively. Next, we bound the average distance of the shortest detour via landmarks. In what follows, we will assume $K(N)=o(N)$ and $K(N)\to\infty$ as $N\to\infty$, implying that the number of landmarks is chosen to be not too large compared to $N$.
Theorem 2.
The average distance of the shortest detour via landmarks, denoted by $\bar{l}$, is bounded above as (2) where $\gamma\approx 0.5772$ is the Euler’s constant.
The proof of Theorem 2 is in Appendix A.2. Next, we apply the results to some well-known models for random graphs.
(4) $\bar{l}_{\textrm{BA}}\leq\dfrac{\log N-\log(\left\langle h^{2}\rangle_{Q}K(N)\right)+\log\log N+\log\log\log N+\log\left[2\log(m/2)/m\right]}{\log\log N+\log(m/2)}+\frac{1}{2}.$
2.4. Erdős-Rényi Model
The Erdős-Rényi (ER) model (Erdős et al., 1960) is a classical random graph in which every node pair is connected with a common probability. For ER graphs, model parameters $\beta$ and $h$ can be set as $\beta=\langle k\rangle N$ and $h\equiv\langle k\rangle$ respectively, where $\langle k\rangle$ denotes the mean degree of nodes (Boguná and Pastor-Satorras, 2003). From (3), we have $q_{ij}=\langle k\rangle/N$, i.e., the probability of edge formation between any node pair is constant. The node degree follows the Poisson distribution with mean $\langle k\rangle$ for large $N$.
We apply our analysis to ER graphs as follows. By plugging in the model parameters to (2), the average distance of the shortest detour via landmarks in ER graphs, denoted by $\bar{l}_{\textrm{ER}}$, is bounded above as
(5) $\bar{l}_{\textrm{ER}}\leq\dfrac{2\log N-\log K(N)}{\log\langle k\rangle}$
asymptotically in $N$. Meanwhile, the average inter-node distance in ER graphs denoted by $\bar{l}^{*}_{\textrm{ER}}$, is given by (Boguná and Pastor-Satorras, 2003; Fronczak et al., 2004)
(6) $\bar{l}^{*}_{\textrm{ER}}=\dfrac{\log N}{\log\langle k\rangle}$
By comparing (5) and (6), we observe that the detour via landmarks incurs the overhead of at most factor 2. This is because, the nodes in ER graphs appear homogeneous, and thus the path length to and from landmarks are on average similar to the inter-node distance. Thus, the average distance of a detour will be twice the direct distance. However, (5) implies that the minimum detour distance can be reduced by using multiple $(K(N)>1)$ landmarks. The reduction can be substantial, e.g., if $K(N)=N^{1-\varepsilon}$ for some $\varepsilon\in(0,1)$, we have, from (5),
$\bar{l}_{ER}\leq(1+\varepsilon)\cdot\bar{l}^{*}_{\textrm{ER}}$
For example, selecting $\sqrt{N}$ landmarks guarantees a 1.5-factor approximation of the shortest path distance. By making $\varepsilon$ close to 0, we get arbitrarily close to the shortest path distance.
Discussion. Due to having Poisson distribution, the degrees in ER graphs are highly concentrated on mean $\langle k\rangle$. There seldom are nodes with very large degrees, i.e., most nodes look alike. Thus, the design question should be on how many rather than on what kind of landmarks. We benefit from choosing a large number of landmarks, e.g., $K(N)=N^{1-\varepsilon}$. However, there is a trade-off: the computational overhead of managing $K(N)$-dimensional vector $D(v)$ will be high.
2.5. Barabási-Albert Model
The Barabási-Albert (BA) model (Barabási and Albert, 1999) generates random graphs with preferential attachment. BA graphs are characterized by continuous growth over time: initially there are $m$ nodes, and new nodes arrive to the network over time. Due to preferential attachment, the probability of a newly arriving node connecting to the existing node is proportional to its degree. The probability of an edge in BA graphs is shown to be (Boguná and Pastor-Satorras, 2003)
$q_{ij}=\frac{m}{2}\frac{1}{\sqrt{t_{i}t_{j}}}$
with $h_{i}=1/\sqrt{t_{i}}$ and $\beta=\frac{2}{m}$, where $t_{i}$ is the time of arrival of node $i$. Since the probability of a newly arriving node connecting to node $i$ is proportional to $h_{i}$, the degree of nodes with large $h_{i}$ is likely to be high. For large $N$, the distribution of $h$ is derived as (Fronczak et al., 2004)
(7) $\rho(h)=\frac{2}{N}h^{-3},~{}h\in\left[\frac{1}{\sqrt{N}},1\right].$
By applying $\rho(\cdot)$ to (2), we bound the average path length with landmarks denoted by $\bar{l}_{\textrm{BA}}$ as (4). Importantly, unlike ER graphs, there exists a landmark selection strategy which achieves the asymptotically optimal distance, despite using a small number of landmarks relative to the network size.
Theorem 3.
Suppose $K(N)$ landmarks are randomly selected from the nodes with top-$(\log N)\cdot K(N)$ largest values of $h$. Assume $m=O(1)$. The average distance of the shortest detour via landmarks, denoted by $\bar{l}_{\textrm{BA}}$, is bounded as
(8) $\bar{l}_{\textrm{BA}}\leq\dfrac{\log N-\log\log K(N)+2\log\log N}{\log\log N}$
asymptotically in $N$.
The proof of Theorem 3 is in Appendix A.3. We claim the optimality of the strategy in Theorem 3 in the following sense. Since $K(N)$ is slowly increasing in $N$, e.g., $K(N)=O(\log N)$, a feasible strategy for Theorem 3 is to choose $O(\log N)$ landmarks from top–$O(\log^{2}N)$ degree nodes. For such $K(N)$, the numerator of the RHS of (8) is $\approx\log N$ for large $N$. Meanwhile, the average length of shortest paths in BA graphs is given by (Cohen and Havlin, 2003)
(9) $\bar{l}^{*}_{\textrm{BA}}=\dfrac{\log N}{\log\log N}$
By comparing (8) and (9), we conclude that $\bar{l}_{\textrm{BA}}$ is asymptotically equal to $\bar{l}^{*}_{\textrm{BA}}$. This implies that the shortest detour via landmarks under our strategy has on average the same length as the shortest path in the asymptotic sense.
Discussion. The node degree in BA graphs is known to follow power-law distribution which causes the emergence of hubs. Inter-node distances can be drastically reduced by hubs, known as small-world phenomenon (Barabási and Albert, 1999). We showed that the shortest path is indeed well-approximated by the detour via landmarks with sufficiently high degrees. Importantly, it is achieved with a small number of landmarks, e.g., let $K(N)=O(\log N)$ in (8).
Summary of Analysis. ER and BA models represent two contrasting cases of degree distributions. The degree distribution of ER graph is highly concentrated, i.e., the variation of node degrees is small, or hubs are absent. By contrast, the degree of BA graphs follows power-law distribution, i.e., the variation of node degree is large, and hubs are present. Our analysis shows that $(1+\epsilon)$-factor approximation (ER) and asymptotic optimality (BA) are achievable by detour via landmarks. The derived bounds are numerically verified via simulation in Appendix B.
[…]
2.6. Design Insights from Theory
The key design questions we would like to address are: what kind of, and how many, landmarks should be selected?
Landmark Selection and Graph Clustering. Theorem 3 states that, one should choose landmarks with sufficiently large $h$, e.g., high-degree nodes. However, such high-degree nodes may not always provide short paths for real-world graphs. For example, suppose all the hubs are located at one end of the network. The nodes at the other end of the network need to make a long detour via landmarks, even in order to reach nodes in local neighborhoods.
In order to better capture local graph structure and evenly distribute landmarks over the graph, we propose to partition the graph into clusters such that the nodes within a cluster tend to be densely connected, i.e., close to one another. Next, we propose to pick the node with the highest degree within each cluster as the landmark, as suggested by the analysis. Each landmark represents the associated cluster, and the distance between nodes can be captured by distance vectors associated with the cluster landmarks. We empirically find that the combination of landmark selection and clustering yields improved results.
Number of Landmarks. Although large $K(N)$ appears preferable, our analysis show that $K(N)$ does not drastically reduce distances, unless $K(N)$ is very large, e.g., $N^{1-\epsilon}$. A large number of landmarks can hamper scalability. Thus, we use only a moderate number of landmarks as suggested by the analysis. Indeed, we empirically found that setting $K(N)=O(\log N)$ suffices to yield good results.
3. Proposed Method
In this section, Hierarchical Position embedding with Landmarks and Clustering (HPLC) is described. The main method combining landmarks and clustering is explained in Sec. 3.1. Additional optimization methods leveraging landmarks and clusters are introduced, which are membership encoding (Sec. 3.2) and cluster-group encoding (Sec. 3.3). An overview of HPLC is depicted in Fig. 1.
3.1. Graph Clustering and Landmark Selection
Graph clustering is a task of partitioning the vertex set into disjoint subsets called clusters, where the nodes within a cluster tend to be densely connected relative to those between clusters (Schaeffer, 2007). We will use FluidC graph clustering algorithm proposed in (Parés et al., 2017) as follows. Suppose we want to partition $G$ to $K$ clusters. FluidC initially selects $K$ random central nodes, assigns each node to clusters $C_{1},\cdots,C_{K}$, and iteratively assigns nodes to clusters according to the following rule. Node $u$ is assigned to cluster $P_{k^{*}(u)}$ such that
(10) $k^{*}(u)=\underset{k=1,\ldots,K}{\arg\!\max}~{}\frac{\left|\left\{u,\mathcal{N}(u)\right\}\cap C_{k}\right|}{|C_{k}|}$
where $\mathcal{N}(u)$ denotes the neighbors of $u$. This assignment rule prefers community candidates of small sizes (denominator), which results in well-balanced cluster sizes. The rule also prefers communities already containing many neighbors of $u$ (numerator), which makes clusters locally dense. As a result, FluidC generates densely-connected cohesive clusters of relatively even sizes. In addition, FluidC has complexity $O(|E|)$, and thus is highly scalable (Parés et al., 2017). Importantly, the number of clusters $K$ can be specified beforehand, which is crucial because $K$ is a hyperparameter in our method.
For each cluster, we select the node with the highest degree as the landmark, where $\lambda_{k}$ denotes the landmark of cluster $C_{k}$. For node $v$, we compute the distances to landmarks $\lambda_{1},\cdots,\lambda_{K}$ to yield distance vector $D(v)$. Since $G$ may be disconnected, we extend the definition of distance $d(v,\lambda)$ such that, in case there exists no path from $v$ to $\lambda$, $d(v,\lambda)$ is set to $\bar{d}+1$ where $\bar{d}$ is the maximum diameter among the connected components. As stated earlier, we set number of clusters $K=\eta\log N$ where integer $\eta$ is a hyperparameter.
Effect of Clustering on Landmarks. The proposed method first performs clustering, and then selects landmarks based on degree centrality. The analysis in Sec. 2 showed that the nodes with sufficiently high degree centrality should be chosen as landmarks. The question is whether the landmarks chosen after clustering will have sufficiently high degrees.
A supporting argument for incorporating clustering into analysis can be made for BA graphs. Theorem 3 states that, choosing $K(N)=O(\log N)$ landmarks from top-$(\log N)\cdot K(N)=O(\log^{2}N)$ nodes in degrees is optimal. We show that, $\log N$ landmarks chosen after FluidC clustering are indeed within top-$(\log N)^{2}$ degree centrality through simulation. Table 1 compares the rank of degrees of landmarks selected after FluidC clustering versus the rank of top-$(\log N)^{2}$ degree nodes in BA graphs. We observe that the landmarks selected after clustering have sufficiently large degrees for optimality, i.e., all of their degrees are within top-$(\log N)^{2}$.
The result can be explained as follows. Power-law graphs are known to have modules which are densely connected subgraphs and are sparsely connected to each other (Goh et al., 2006). A proper clustering algorithm is expected to detect modules. Moreover, each cluster contains a relatively large number of nodes, because the number of clusters is relatively small $(O(\log N))$. Thus, each cluster is likely to contain nodes with sufficiently high degrees which are, according to Theorem 3, good candidates for landmarks.
One can expect similar effects from FluidC clustering for real-world graphs. Given the small number $(O(\log N))$ of clusters as input, the algorithm will detect locally dense clusters resulting in landmarks which have high degrees and tend to be evenly distributed over the graph. Thus, landmark selection after clustering can be a good heuristic motivated by the theory.
Network size $N$ rank: cluster landmarks rank: top-$(\log N)^{2}$ Fraction of landmarks within rank top-$(\log N)^{2}$ 500 5.02% 6.38% 100% 1000 3.02% 4.37% 100% 2000 1.76% 2.85% 100% 5000 0.90% 1.48% 100%
Table 1. Rank of degrees of landmark nodes in BA graphs. For example, if $N=500$, all the landmarks selected from clustering are within the top 5.02%-degree nodes.
3.2. Membership Encoding with Graph Laplacian
For the nodes within the same cluster, we augment the nodes’ embeddings with information identifying that they are members of the same community, which we call membership. The membership is extracted from landmarks, and is based on the relative positional information among landmarks. Thus, not only the nodes within the same cluster have the same membership, but also the nodes of neighboring clusters have “similar” membership. The membership of a node is encoded into a membership vector (MV) which is combined with DV in computing the node embeddings.
We consider the graph consisting only of landmark nodes, and use graph Laplacian (Belkin and Niyogi, 2003) to encode their relative positions as follows. The landmark graph uses the edge weight between landmarks $u$ and $v$ which is set to $e_{uv}=\exp(-d(u,v)^{2}/T)$ where $T$ denotes normalizing parameter of heat kernel. Let $\hat{A}\in\mathbb{R}^{K\times K}$ denote the weighted adjacency matrix. The normalized graph Laplacian is given by $L=I-\Delta^{-\frac{1}{2}}\hat{{A}}\Delta^{-\frac{1}{2}}$ where degree matrix $\Delta$ is given by $\Delta_{ii}=\sum_{j}\hat{A}_{ij}$. The eigenvectors of $L$ are used as MVs, i.e., for node $v\in C_{k}$, the MV of node $v$, denoted by $M(v)$, is the eigenvector of $L$ associated with landmark $\lambda_{k}$. MV provides positional information in addition to DV, using the graph of upper hierarchy, i.e., landmarks. We use a random flipping of the signs of eigenvectors to resolve ambiguity (Dwivedi et al., 2020). A problem with Laplacian encoding is the time complexity: one needs the eigendecomposition whose complexity is cubic in network size. Thus, previous approaches had limitations, i.e., used only a subset of eigenspectrum (Dwivedi et al., 2020; Kreuzer et al., 2021). However, in our method, the landmark graph has small size, i.e., $K=O(\log N)$. Thus, we can exploit the full eigenspectrum, enabling more accurate representation of landmark graphs.
3.3. Cluster-group Encoding
We propose cluster-group encoding as an additional model optimization as follows. Several neighboring clusters are further grouped into a macro-cluster. The embeddings of nodes in a macro-cluster use an encoder specific to that macro-cluster. The motivation is that, using a separate encoder per local region may facilitate capturing attributes specific to local structures or latent features of the community, which is important for link prediction.
For cluster $C_{k}$, let $\Phi_{i(k)}$ denote the macro-cluster which contains $C_{k}$ where $i(k)$ denote the index of the macro-cluster. The node embedding $z_{v}$ for node $v$ is computed as
$z_{v}=f_{i(k)}\left(x_{v}\oplus\phi\left(M(v)\oplus D(v)\right)\right),\quad v\in C_{k}\subset\Phi_{i(k)},$
where $x_{v}$ is the input node feature, $M(v)$ and $D(v)$ are MV and DV, $\phi$ is membership/distance encoder, and $\oplus$ denotes concatenation. $f_{i(\cdot)}$ denotes the encoder specific to macro-cluster $\Phi_{i(\cdot)}$. MLPs are used for $\phi(\cdot)$ and $\Phi_{i(\cdot)}$. The number of macro-cluster, denoted by $R$, is set such that $K$ is a multiple of $R$. Each macro-cluster contains $K/R$ clusters. Cluster-group encoding requires $R$ encoders, one per macro-cluster. To limit the model size, we empirically set $R=\min(15,\lfloor K/\eta\rfloor)$. For ease of implementation, $G$ is first partitioned into $R$ macro-clusters, then each macro-cluster is partitioned into $K/R$ clusters. This is simpler than, say, partitioning $G$ into $K$ (smaller) clusters and grouping them back to macro-clusters.
Next, $z_{v}$ is input to GNN whose $l$-th layer output $h^{l}_{v}$ is given by
$h^{l}_{v}=\textrm{GNN}(h^{l-1}_{v},\mathbf{A}),\quad\textrm{for }l=1,2,\ldots,L$
where $h^{0}_{v}=z_{v}$. HPLC is a plug-in method and thus can be combined with different types of GNNs. The related study is provided in Sec. 7.2. Finally, we concatenate node-pair embeddings and compute scores by $y_{v,u}=\sigma(h^{L}_{v}\oplus h^{L}_{u})$ where $y_{v,u}$ denotes the link prediction score between node $v$ and $u$, and $\sigma(\cdot)$ is an MLP.
4. Property of HPLC as Node embedding
In (Srinivasan and Ribeiro, 2020), the (positional) node embeddings are formally defined as:
Definition 0.
The node embeddings of a graph with adjacency matrix $\mathbf{A}$ and input attributes $X$ are defined as joint samples of random variables $(Z_{i})_{i\in V}|\mathbf{A},X\sim p(\cdot|\mathbf{A},X),$ $Z_{i}\in\mathbb{R}^{d}$, $d\geq 1$, where $p(\cdot|\mathbf{A},X)$ is a $\mathcal{G}$-equivariant probability distribution on $\mathbf{A}$ and $X$, that is, $\pi(p(\cdot|\mathbf{A},X))=p(\cdot|\pi(\mathbf{A}),\pi(X))$ for any permutation $\pi(\cdot).$
It is argued in (Srinivasan and Ribeiro, 2020) that positional node embeddings are good at link prediction tasks: the positional embeddings preserve the relative positions of nodes, which enables differentiating isomorphic nodes according to their positions and identifying the closeness between nodes. We claim that HPLC qualifies as positional node embeddings according to Definition 1 as follows.
In HPLC, landmark-based distance function, eigenvectors of graph Laplacian with random flipping, and graph clustering method are $\mathcal{G}$-equivariant functions of $\mathbf{A}$ ignoring the node features $X$. The landmark-based distance vectors and its functions clearly satisfy the equivariance property. The eigenvectors of graph Laplacian are permutationally equivariant under the node permutation, i.e., switching of corresponding rows/columns of adjacency matrix. Also, the Laplacian eigenmap can be regarded as a $\mathcal{G}$-equivariant function in the sense of expectation, if it is combined with random sign flipping of eigenvectors (Srinivasan and Ribeiro, 2020). In case of multiplicity of eigenvalues, we can slightly perturb the edge weights of landmark graphs to obtain simple eigenvalues (Poignard et al., 2018). The graph clustering in HPLC is a randomized method, because initially $K$ central nodes are selected at random. Thus, embedding output $Z$ of HPLC is a function of $\mathbf{A},X$, and random noise, which proves our claim that HPLC qualifies as a positional node embedding.
Moreover, HPLC is more expressive than standard message-passing GNNs as follows. HPLC is trained to predict the link between a node pair based on their positional embeddings. The embeddings are learned over the joint distribution of distances from node pairs to common landmarks which are spread out globally over the graph. By contrast, traditional GNNs learn the embeddings based on the marginal distributions of local neighborhoods of node pairs. By a similar argument to Sec. 5.2 of (You et al., 2019), HPLC induces higher mutual information between its embeddings and the target similarity (likelihood of link formation) of node pairs based on positions, and thus has higher expressive power than standard GNNs.
5. Complexity Analysis
5.1. Time Complexity
The time complexity of HPLC is incurred mainly from computing DVs and MVs. We first consider the complexity of computing $D(v)$ for all $v\in V$. There are $\eta\log N$ landmarks, and for each landmark, computing the distances from all $v\in V$ to the landmark requires $O(|E|+N\log N)$ using Fibonacci heap. Thus, the overall complexity is $O(|E|\log N+N(\log N)^{2})$. Next, we consider the complexity of computing $M(v)$. We compute Laplacian eigenvectors of landmark graph, where its time complexity is $O(K^{3})=O((\log N)^{3})$. Importantly, the computation of $D(v)$ and $M(v)$ can be done during the preprocessing stage.
5.2. Space Complexity
The space complexity of HPLC is mainly from the GNN models for computing the node embeddings. Additional space complexity of HPLC is from the membership/distance encoder $\phi(\cdot)$ which is $O\left((F+R)H_{\textrm{in}}\right)$, and cluster-group encoding which is $O\left(H_{\textrm{in}}H_{\textrm{out}}R\right)$ respectively, where $F$ denotes the node feature dimension, $R=\min(15,\lfloor K/\eta\rfloor)$ denotes the number of macro-clusters, $H_{\textrm{in}}$ denotes the hidden dimension of $\phi(\cdot)$, and $H_{\textrm{out}}$ denotes the hidden dimension of cluster-group encoders.
Overall, the time and space complexity of HPLC is low, and our experiments shows that HPLC handles large or dense graphs well. We also measured the actual resource usage, and the related results are provided in Appendix D.
Baselines Avg. (H.M.) CITATION2 COLLAB DDI PubMed Cora Citeseer Facebook Adamic Adar 65.74 (50.65) 76.12 $\pm$ 0.00 53.00 $\pm$ 0.00 18.61 $\pm$ 0.00 66.89 $\pm$ 0.00 77.22 $\pm$ 0.00 68.94 $\pm$ 0.00 99.41 $\pm$ 0.00 MF 54.06 (42.65) 53.08 $\pm$ 4.19 38.74 $\pm$ 0.30 17.92 $\pm$ 3.57 58.18 $\pm$ 0.01 51.14 $\pm$ 0.01 50.54 $\pm$ 0.01 98.80 $\pm$ 0.00 Node2Vec 64.01 (51.17) 53.47 $\pm$ 0.12 41.36 $\pm$ 0.69 21.95 $\pm$ 1.58 80.32 $\pm$ 0.29 84.49 $\pm$ 0.49 80.00 $\pm$ 0.68 86.49 $\pm$ 4.32 GCN (GAE) 76.35 (66.67) 84.74 $\pm$ 0.21 44.14 $\pm$ 1.45 37.07 $\pm$ 5.07 95.80 $\pm$ 0.13 88.68 $\pm$ 0.40 85.35 $\pm$ 0.60 98.66 $\pm$ 0.04 GCN (MLP) 76.48 (67.53) 84.79 $\pm$ 0.24 44.29 $\pm$ 1.88 39.31 $\pm$ 4.87 95.83 $\pm$ 0.80 90.25 $\pm$ 0.53 81.47 $\pm$ 1.40 99.43 $\pm$ 0.02 GraphSAGE 78.51 (71.48) 82.64 $\pm$ 0.01 48.62 $\pm$ 0.87 44.82 $\pm$ 7.32 96.58 $\pm$ 0.11 90.24 $\pm$ 0.34 87.37 $\pm$ 1.39 99.29 $\pm$ 0.01 GAT - - 44.14 $\pm$ 5.95 29.53 $\pm$ 5.58 85.55 $\pm$ 0.23 82.59 $\pm$ 0.14 87.29 $\pm$ 0.11 99.37 $\pm$ 0.00 JKNet - - 48.84 $\pm$ 0.83 57.98 $\pm$ 7.68 96.58 $\pm$ 0.23 89.05 $\pm$ 0.67 88.58 $\pm$ 1.78 99.43 $\pm$ 0.02 P-GNN - - - 1.14 $\pm$ 0.25 87.22 $\pm$ 0.51 85.92 $\pm$ 0.33 90.25 $\pm$ 0.42 93.13 $\pm$ 0.21 GTrans+LPE - - 11.19 $\pm$ 0.42 9.22 $\pm$ 0.20 81.15 $\pm$ 0.12 79.31 $\pm$ 0.09 77.49 $\pm$ 0.02 99.27 $\pm$ 0.00 GCN+LPE 74.60 (67.00) 84.85 $\pm$ 0.35 49.75 $\pm$ 1.35 38.18 $\pm$ 7.62 95.50 $\pm$ 0.13 76.46 $\pm$ 0.15 78.29 $\pm$ 0.21 99.17 $\pm$ 0.00 GCN+DE 72.48 (60.04) 60.30 $\pm$ 0.61 53.44 $\pm$ 0.29 26.63 $\pm$ 6.82 95.42 $\pm$ 0.08 89.51 $\pm$ 0.12 86.49 $\pm$ 0.11 99.38 $\pm$ 0.02 GCN+LRGA 78.42 (74.47) 65.05 $\pm$ 0.22 52.21 $\pm$ 0.72 62.30 $\pm$ 9.12 93.53 $\pm$ 0.25 88.83 $\pm$ 0.01 87.59 $\pm$ 0.03 99.42 $\pm$ 0.05 SEAL 77.08 (62.88) 85.26 $\pm$ 0.98 53.72 $\pm$ 0.95 26.25 $\pm$ 8.00 95.86 $\pm$ 0.28 92.55 $\pm$ 0.50 85.82 $\pm$ 0.44 99.60 $\pm$ 0.02 NBF-net - - - 4.03 $\pm$ 1.32 97.30 $\pm$ 0.45 94.12 $\pm$ 0.17 92.30 $\pm$ 0.23 99.42 $\pm$ 0.04 PEG-DW+ 81.67 (75.42) 86.03 $\pm$ 0.53 53.70 $\pm$ 1.18 47.88 $\pm$ 4.56 97.21 $\pm$ 0.18 93.12 $\pm$ 0.12 94.18 $\pm$ 0.18 99.57 $\pm$ 0.05 HPLC 85.77 (82.39) 86.15 $\pm$ 0.48 56.04 $\pm$ 0.28 70.03 $\pm$ 7.02 97.38 $\pm$ 0.34 94.95 $\pm$ 0.18 96.15 $\pm$ 0.19 99.69 $\pm$ 0.00
Table 2. Link prediction results on various datasets. All baselines and our method were evaluated for 10 repetitions. Bold denotes the best performance, and Italic indicates the second best performance. We used a single NVIDIA RTX 3090 with 24GB memory on all datasets except CITATION2 and A100 GPU with 40GB memory on CITATION2. - indicates ‘out-of-memory’ (OOM). Some baselines suffered from OOM on large graphs due to the high memory usage from storing a large number of shortest paths, attention weights, or aggregation of hidden embedding vectors, etc. Similar OOM results as well as poor performance of those baselines were reported in (Wang et al., 2022) and (Zhao et al., 2022). For SEAL and GCN+DE, we trained 2% of training data and evaluated 1% of both validation and test set respectively on CITATION2. We trained 15% of training data but evaluated all of the validation and test sets on COLLAB. Both implementations followed the guideline on the official GitHub of SEAL-OGB. ‘Avg.’ denotes the average of performance metrics, and ‘H.M’ indicates their harmonic mean. (-) in ‘Avg. (H.M.)’ means that we do not report the average and harmonic mean due to OOM.
6. Experiments
6.1. Experimental setting
Datasets. Experiments were conducted on 7 datasets widely used for evaluating link prediction. For experiments on small graphs, we used PubMed, Cora, Citeseer, and Facebook. For experiments on dense or large graphs, we chose DDI, COLLAB, and CITATION2 provided by OGB (Hu et al., 2020). Detailed statistics and evaluation metrics of the datasets are provided in Table 9 in Appendix E.
Baseline models. We compared HPLC with Adamic Adar (AA) (Adamic and Adar, 2003), Matrix Factorization (MF) (Koren et al., 2009), Node2Vec (Grover and Leskovec, 2016), GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), P-GNN (You et al., 2019), NBF-net (Zhu et al., 2021), and plug-in type approaches like JKNet (Xu et al., 2018), SEAL (Zhang and Chen, 2018), GCN+DE (Li et al., 2020), GCN+LPE (Dwivedi et al., 2020), GCN+LRGA (Puny et al., 2021), Graph Transformer+LPE (Dwivedi and Bresson, 2021), and PEG-DW+ (Wang et al., 2022). All methods except for AA and GAE are computed by the same decoder, which is a 2-layer MLP. For a fair comparison, we use GCN in most plug-in type approaches: SEAL, GCN+DE, GAE, JKNet, GCN+LRGA, GCN+LPE, and HPLC.
Evaluation metrics. Link prediction was evaluated based on the ranking performance of positive edges in the test data over negative ones. For COLLAB and DDI, we ranked all positive and negative edges in the test data, and computed the ratio of positive edges which are ranked in top-$k$. We did not utilize validation edges for computing node embeddings when we predicted test edges on COLLAB. In CITATION2, we computed all positive and negative edges, and calculated the reverse of the mean rank of positive edges. Due to high complexity when evaluating SEAL, we only trained 2% of training set edges and evaluated 1% of validation and test set edges respectively, as recommended in the official GitHub of SEAL. The metrics for DDI, COLLAB and CITATION2 are chosen according to the official OGB settings  (Hu et al., 2020). For Cora, Citeseer, PubMed, and Facebook, Area Under ROC Curve (AUC) is used as the metric similar to prior works (Kipf and Welling, 2017; Wang et al., 2022). If applicable, we calculated the average and harmonic mean (HM) of the measurements. HM penalizes the model for very low scores, thus is a useful indicator of robustness.
Hyperparameters. We used GCN as our base GNN encoder. MLP is used in decoders, except GAE. None of the baseline methods used edge weights. During training, negative edges are randomly sampled at a ratio of 1:1 with positive edges. The details of hyperparameters are provided in Appendix F.
6.2. Results
Experimental results are summarized in Table 2. HPLC outperformed the baselines on most datasets. HPLC achieved large performance gains over GAE combined with GCN on all datasets, which are 88.9% on DDI, 27.0% on COLLAB, 12.7% on Citeseer, 7.1% on Cora, and 1.4% on CITATION2. HPLC showed superior performance over SEAL, achieving gains of 167% on DDI, and 12.0% on Citeseer. We compare HPLC with other distance-based methods. Compared to GCN+DE which encodes distances from a target node set whose representations are to be learned, or to P-GNN which uses distances to random anchor sets, HPLC achieved higher performance gains by a large margin. HPLC outperformed other positional encoding methods, e.g., GCN+LPE, Graph Transformer+LPE, and PEG-DW+. The results show that the proposed landmark-based representation can be effective for estimating the positional information of nodes.
SEAL and NBF-net performed poorly on DDI which is a highly dense graph. Since the nodes of DDI have a large number of neighbors, the enclosing subgraphs are both very dense and large, and the models struggle with learning the representations of local structures or paths between nodes. By contrast, HPLC achieved the best performance on DDI, demonstrating its effectiveness on densely connected graphs.
Finally, we computed the average and harmonic means of performance measurements except for the methods with OOM problems. Although the averages are taken over heterogeneous metrics and thus do not represent specific performance metrics, they are presented for comparison purposes. In summary, HPLC achieved the best average and harmonic mean of performance measurements, demonstrating both its effectiveness and robustness.
7. Ablation study
In this section, we provide the ablation study. Additional ablation studies on graph clustering algorithms and node centrality are provided in Appendix C.
7.1. Model Components
Table 3 shows the performances in the ablation analysis for the model components. The components denoted by“DV”, “CE” and “MV” columns in Table 3 indicate the usage of distance vector, cluster-group encoding and membership vector, respectively. The results show that all of the hierarchical components, i.e., “DV”, “CE” and “MV”, indeed contribute to the performance improvement.
DV CE MV COLLAB DDI PubMed Cora Citeseer
44.29 $\pm$ 1.88 39.31 $\pm$ 4.87 95.83 $\pm$ 0.80 90.25 $\pm$ 0.53 81.43 $\pm$ 0.02
52.64 $\pm$ 0.39 54.00 $\pm$ 8.90 95.93 $\pm$ 0.19 92.21 $\pm$ 0.13 95.54 $\pm$ 0.16
44.33 $\pm$ 0.24 41.33 $\pm$ 5.82 95.92 $\pm$ 0.15 91.82 $\pm$ 0.18 94.87 $\pm$ 0.12
53.31 $\pm$ 0.62 46.86 $\pm$ 9.91 95.97 $\pm$ 0.13 92.32 $\pm$ 0.25 95.13 $\pm$ 0.18
55.56 $\pm$ 0.37 68.75 $\pm$ 7.43 96.67 $\pm$ 0.18 93.94 $\pm$ 0.21 95.83 $\pm$ 0.15
56.04 $\pm$ 0.28 70.03 $\pm$ 7.02 97.38 $\pm$ 0.34 94.95 $\pm$ 0.18 96.15 $\pm$ 0.19
Table 3. Ablation study on model components.
7.2. Combination with various GNNs
HPLC can be combined with different GNN encoders. We experimented the combination with three types of widely-used GNN encoders. Table 4 shows that, HPLC enhances the performance of various types of GNNs.
Dataset w/ GraphSAGE w/ GAT w/ GCN
PubMed 96.63 ($+0.05$) 93.18 ($+7.63$) 97.38 ($+1.55$)
Cora 96.18 ($+5.94$) 91.93 ($+9.34$) 94.95 ($+4.0$)
Citeseer 94.97 ($+7.60$) 94.60 ($+7.31$) 96.15 ($+14.68$)
Facebook 99.48 ($+0.19$) 99.40 ($+0.03$) 99.69 ($+0.26$)
Table 4. Ablation study on GNN types. $+$ denotes the performance gain over the default GNN encoder without HPLC.
8. Related Work
Link Prediction. Earlier methods for link prediction used heuristics (Adamic and Adar, 2003; Page et al., 1998; Zhou et al., 2009) based on manually designed formulas. GNNs were subsequently applied to the task, e.g., GAE (Kipf and Welling, 2016) is a graph auto-encoder which reconstructs adjacency matrices combined with GNNs, but cannot distinguish isomorphic nodes. SEAL (Zhang and Chen, 2018) is proposed as structural link representation by extracting enclosing subgraphs and learning structural patterns of those subgraphs. The authors demonstrated that higher-order heuristics can be approximately represented by lower-order enclosing subgraphs thanks to $\gamma$-decaying heuristic. LGLP (Cai et al., 2021) used the graph transformation prior to GNN layers for link prediction, and Walk Pooling (Pan et al., 2022) proposed to learn subgraph structure based on random walks. However, the aforementioned methods need to extract enclosing subgraphs of edges and compute their node embeddings on the fly. CFLP (Zhao et al., 2022) is a counterfactual learning framework for link prediction to learn causal relationships between nodes. However, its time complexity is $O(N^{2})$ for finding counterfactual links with nearest neighbors.
Distance- and Position-based Methods. P-GNN (You et al., 2019) proposed position-aware GNN to inject positional information based on distances into node embeddings. P-GNN focuses on realizing Bourgain’s embedding (Bourgain, 1985) guided by Linial’s method (Linial et al., 1995) and performs message computation and aggregation based on distances to random subset of nodes. By contrast, we judiciously select representative nodes in combination with graph clustering and use the associated distances. Laplacian positional encodings (Belkin and Niyogi, 2003; Dwivedi et al., 2020) use eigenvectors of graph Laplacian as positional embeddings in which positional features of nearby nodes are encoded to be similar to one another. Graph transformer was combined with positional encoding learned from Laplacian spectrum  (Dwivedi and Bresson, 2021; Kreuzer et al., 2021). However, transformers with full attention have high computational complexity and do not scale well for link predictions in large graphs. Distance encoding (DE) as node labels was proposed and its expressive power was analyzed in (Li et al., 2020). In (Zhang et al., 2021), the authors analyzed the effects of various node labeling tricks using distances. However, these two methods do not utilize distances as positional information.
Networks with landmarks. Algorithms augmented with landmarks have actively been explored for large networks, where the focus is mainly on estimating the inter-node distances (Ng and Zhang, 2002; Tang and Crovella, 2003; Dabek et al., 2004; Zhao and Zheng, 2010) or computing shortest paths (Goldberg and Harrelson, 2005; Tretyakov et al., 2011; Akiba et al., 2013; Qiao et al., 2012). An approximation theory on inter-node distances using embeddings derived from landmarks is proposed in (Kleinberg et al., 2004) which, however, based on randomly selected landmarks, whereas we analyze detour distances under a judicious selection strategy. In (Potamias et al., 2009), vectors of distances to landmarks are used to estimate inter-node distances, and landmark selection strategies based on various node centralities are proposed. However, the work did not provide theoretical analysis on the distances achievable under detours via landmarks. The aforementioned works do not consider landmark algorithms in relation to link prediction tasks.
9. Conclusion
We proposed a hierarchical positional embedding method using landmarks and graph clustering for link prediction. We provided a theoretical analysis of the average distances of detours via landmarks for well-known random graphs. From the analysis, we gleaned design insights on the type and number of landmarks to be selected and proposed HPLC which effectively infuses positional information using $O(\log N)$-landmarks for the link prediction on real-world graphs. Experiments demonstrated that HPLC achieves state-of-the-art performance and has better scalability as compared to existing methods on various graph datasets of diverse sizes and densities. In the future, we plan to analyze the landmark strategies for various types of random networks and extend HPLC to other graph-related tasks such as graph classification, graph generation, etc.
10. Acknowledgement
This research was supported by the MSIT (Ministry of Science and ICT), Korea, under the ICT Creative Consilience program (IITP-2024-2020-0-01819) supervised by the IITP (Institute for Information & communications Technology Planning & Evaluation), and by the National Research Foundation of Korea (NRF) Grant through MSIT under Grant No. 2022R1A5A1027646 and No. 2021R1A2C1007215.
Appendix A Proofs
A.1. Proof of Theorem 1.
The framework from (Fronczak et al., 2004) is used for the proof, and their technique is briefly described as follows. We first state a key lemma from (Fronczak et al., 2004):
Lemma 0.
If $A_{1},\cdots,A_{n}$ are mutually independent events and their probabilities fulfill relations $\forall_{i}P(A_{i})\leq\varepsilon$,
$P\left(\bigcup_{k=1}^{n}A_{k}\right)=1-\exp\left(-\sum_{k=1}^{n}P(A_{k})\right)-R$
where $0\leq R<\sum_{j=0}^{n+1}\left(n\varepsilon\right)^{j}/j!-\left(1+\varepsilon\right)^{n}$. Thus, $R$ vanishes in the limit $n\to\infty$.
Consider node $i,j\in V$ and landmark $\lambda\in V$. Let $p^{\lambda}_{ij}(s)$ denote the probability that the length of paths between $i$ and $j$ via $\lambda$ is at most $s$ for $s=1,2,\cdots$. $p^{\lambda}_{ij}(s)$ is equivalent to the probability that there exists at least one walk (i.e., revisiting a node is allowed) of length $s$ from $i$ to $j$ via $\lambda$. The probability of the existence of a walk of length $s$ in a specific node sequence $i\to v_{1}\to\ldots\to v_{s-1}\to j$ is given by
$q_{iv_{1}}q_{v_{1}v_{2}}\cdots q_{v_{s-1}j}$
where $q_{ij}$ is the edge probability as defined in (3).
We claim that $p^{\lambda}_{ij}(s)$ can be expressed as
(11) $p^{\lambda}_{ij}(s)=1-\exp\left[-\left\{\sum_{v_{1}=1}^{N}\cdots\sum_{v_{s-1}=1}^{N}q_{iv_{1}}q_{v_{1}v_{2}}\cdots q_{v_{s-1}j}-\sum_{\begin{subarray}{c}v_{1}=1\\ v_{1}\neq\lambda\end{subarray}}^{N}\cdots\sum_{\begin{subarray}{c}v_{s-1}=1\\ v_{s-1}\neq\lambda\end{subarray}}^{N}q_{iv_{1}}q_{v_{1}v_{2}}\cdots q_{v_{s-1}j}\right\}\right]$
The first summation in the bracket of (11) counts all possible walks of length $s$ from $i$ to $j$ and sums up their probabilities. The second summation in the bracket of (11) counts all possible walks from $i$ to $j$ but never visits $\lambda$. Thus the subtraction in the bracket counts all the walks from $i$ to $j$ visiting $\lambda$ at least once. Thus, expression (11) is the probability of the existence of walks of length $s$ from $i$ to $j$ via $\lambda$ from Lemma 1, e.g., $A_{k}$ in Lemma 1 corresponds to an event of a walk. The expression is asymptotically accurate in $N$, i.e., although Lemma 1 requires events $A_{k}$ be independent, and the same edge may participate between different $A_{k}$’s and induce correlation, the fraction of such correlations becomes negligible when $s\ll N$, as argued in (Fronczak et al., 2004) [^1] Alternatively, a simple proof can be done by using a counting argument: the ratio of the number of paths not visiting the same node more than once (which are uncorrelated paths) out of all the possible paths of length $s$ from node $i$ to $j$ is close to 1 if $s\ll N$..
Let us evaluate the summations in the bracket of (11). We have
$\sum_{v_{1}=1}^{N}\cdots\sum_{v_{s-1}=1}^{N}q_{iv_{1}}q_{v_{1}v_{2}}\cdots q_{v_{s-1}j}=h_{i}h_{j}\frac{\left(N\langle h^{2}\rangle\right)^{s-1}}{\beta^{s}}$
whereas
$\sum_{\begin{subarray}{c}v_{1}=1\\ v_{1}\neq\lambda\end{subarray}}^{N}\cdots\sum_{\begin{subarray}{c}v_{s-1}=1\\ v_{s-1}\neq\lambda\end{subarray}}^{N}q_{iv_{1}}q_{v_{1}v_{2}}\cdots q_{v_{s-1}j}=h_{i}h_{j}\frac{\left(N\langle h^{2}\rangle-h_{\lambda}^{2}\right)^{s-1}}{\beta^{s}}$
Thus the subtraction in the bracket of (11) is given by
$\frac{h_{i}h_{j}}{\beta^{s}}\left[\left(N\langle h^{2}\rangle\right)^{s-1}-\left(N\langle h^{2}\rangle-h_{\lambda}^{2}\right)^{s-1}\right]$
(12) $=\frac{h_{i}h_{j}}{\beta^{s}}\left(N\langle h^{2}\rangle\right)^{s-1}\left[1-\left(1-\frac{h_{\lambda}^{2}}{N\langle h^{2}\rangle}\right)^{s-1}\right]$
(13) $\approx\frac{h_{i}h_{j}}{\beta^{s}}\left(N\langle h^{2}\rangle\right)^{s-1}(s-1)\frac{h_{\lambda}^{2}}{N\langle h^{2}\rangle}$
(14) $=\frac{h_{i}h_{j}h_{\lambda}^{2}}{\beta N\langle h^{2}\rangle}(s-1)\left(\frac{N\langle h^{2}\rangle}{\beta}\right)^{s-1}$
where approximation (13) can be shown to be valid for large $N$, i.e., $\frac{h^{2}_{\lambda}}{N\langle h^{2}\rangle}\ll 1$ holds, assuming $h$ has a finite second-order moment $\langle h^{2}\rangle$. Let random variable $L_{ij}(\lambda)$ denote the path length from node $i$ to $j$ with visiting landmark $\lambda$. Then we have $p^{\lambda}_{ij}(s)=P(L_{ij}(\lambda)\leq s)$. Let
$F_{\lambda}(s):=P(L_{ij}(\lambda)>s).$
We have that
$F_{\lambda}(s)=1-p^{\lambda}_{ij}(s)=\exp\left[-\frac{h_{i}h_{j}h_{\lambda}^{2}}{\beta N\langle h^{2}\rangle}(s-1)\left(\frac{\langle h^{2}\rangle N}{\beta}\right)^{s-1}\right]$
for $s=1,2,\cdots$.
Consider the minimum distance among the routes via landmarks $\lambda_{k}$, $k=1,\cdots,K(N)$ where the landmarks are chosen i.i.d. from distribution $Q$. Let $L_{ij}$ denote the minimum distance among the routes from $i$ to $j$ via the landmarks. Then
$L_{ij}=\min[L_{ij}(\lambda_{1}),L_{ij}(\lambda_{2}),\cdots,L_{ij}(\lambda_{K(N)})].$
We have that
$P(L_{ij}>s)$
$=P(\min[L_{ij}(\lambda_{1}),L_{ij}(\lambda_{2}),\cdots,L_{ij}(\lambda_{K(N)})]>s)$
$=P(L_{ij}(\lambda_{1})>s,L_{ij}(\lambda_{2})>s,\cdots,L_{ij}(\lambda_{K(N)})>s)$
$=\prod_{k=1}^{K(N)}P(L_{ij}(\lambda_{k})>s)$
$=\exp\left[-\frac{h_{i}h_{j}}{\beta N\langle h^{2}\rangle}\left(\sum_{k=1}^{K(N)}h_{\lambda_{k}}^{2}\right)(s-1)\left(\frac{\langle h^{2}\rangle N}{\beta}\right)^{s-1}\right]$
(15) $=\exp\left[-\frac{h_{i}h_{j}}{\beta N\langle h^{2}\rangle}K(N)\cdot\langle h^{2}\rangle_{Q}\cdot(s-1)\left(\frac{\langle h^{2}\rangle N}{\beta}\right)^{s-1}\right]$
which proves (1), where we have used
$\left(\sum_{k=1}^{K(N)}h_{\lambda_{k}}^{2}\right)$ $=K(N)\cdot\frac{1}{K(N)}\left(\sum_{k=1}^{K(N)}h_{\lambda_{k}}^{2}\right)$
$\approx K(N)\cdot\langle h^{2}\rangle_{Q}$
in (15) for sufficiently large $K(N)$ where $\langle\cdot\rangle_{Q}$ denotes the expectation of the hidden variables of landmarks chosen according to distribution $Q$.
A.2. Proof of Theorem 2.
Let $l_{ij}$ denote the mean length of the shortest detour via landmarks from $i$ to $j$. We have that
$l_{ij}$ $=\sum_{s=1}^{\infty}P(L_{ij}>s)$
$=\sum_{s=0}^{\infty}\exp\left[-\frac{h_{i}h_{j}}{\beta N\langle h^{2}\rangle}\cdot\langle h^{2}\rangle_{Q}K(N)\cdot s\left(\frac{\langle h^{2}\rangle N}{\beta}\right)^{s}\right]$
We utilize the Poisson summation formula (Guinand, 1941):
[…]
where
(17) $f(t)$ $=\exp\left[-atb^{t}\right],$
(18) $a$ $:=\langle h^{2}\rangle_{Q}K(N)\cdot\frac{h_{i}h_{j}}{\beta N\langle h^{2}\rangle},$
(19) $b$ $:=\frac{\langle h^{2}\rangle N}{\beta}$
Firstly we have $f(0)=1$. Next, we evaluate the second term of (16):
[…]
Let $u=te^{t}$, then we have
$dt=\frac{du}{u+e^{W(u)}}$
where $W(\cdot)$ is the Lambert $W$ function which is the inverse of function $te^{t}$ for $t\geq 0$. Thus (20) is equal to
[…]
Since $W(u)\geq 0$ for $u\geq 0$, (20) is bounded above by
[…]
(21) $=-\exp\left(\frac{a}{\log b}\right)\frac{\mathrm{Ei}\left(-\frac{a}{\log b}\right)}{\log b}$
where $\mathrm{Ei}(\cdot)$ denotes the exponential integral. Consider the assumption $K(N)=o(N)$, i.e., the number of landmarks is not too large compared to $N$. Under this assumption, one can verify that $a/\log b$ is at most $o(N)/N$ which tends to 0 as $N\to\infty$, and the exponential term of (21) can be approximated to 1. Thus, (21) reduces to
$-\frac{\mathrm{Ei}\left(-\frac{a}{\log b}\right)}{\log b}$ $=\dfrac{-\gamma-\log a+\log\log b}{\log b}$
where we used
$-\mathrm{Ei}\left(-\frac{a}{\log b}\right)\approx-\gamma-\log\left(\frac{a}{\log b}\right)$
where the error term associated with exponential integral vanishes because $a/\log b$ is small, and $\gamma\approx 0.5772$ is the Euler’s constant. By applying the approximation and plugging in $a$ and $b$ to (21), we get expression
(22) $\dfrac{-\log(h_{i}h_{j})-\log\left(\langle h^{2}\rangle_{Q}K(N)\right)+\log(N\beta\langle h^{2}\rangle)+\log\log\left(\frac{N\langle h^{2}\rangle}{\beta}\right)-\gamma}{\log N+\log\langle h^{2}\rangle-\log\beta}$
for the second term of (16).
Finally, one can show that the last term of (16) vanishes, by using generalized mean value theorem (Fronczak et al., 2004). By averaging (22) over all $i,j\in V$, we obtain (2).
A.3. Proof of Theorem 3
Let $M(N):=(\log N)\cdot K(N)$. The landmarks are selected at random from $M(N)$ nodes with highest values of $h$. This implies that the distribution $Q(\cdot)$ of $h$ values of landmarks is given by the following conditional distribution:
$Q(t)$ $=\rho(t|h\geq\frac{1}{\sqrt{M(N)}})$
(23) $={\rho(t)\cdot\mathbf{1}\bigg{(}t\in\bigg{[}\frac{1}{\sqrt{M(N)}},1\bigg{]}\bigg{)}}\bigg{/}P\bigg{(}h\in\bigg{[}\frac{1}{\sqrt{M(N)}},1\bigg{]}\bigg{)}\bigg{.}$
where $\rho(\cdot)$ is the distribution of $h$ given by (7).
From (23), we have that
$\langle h^{2}\rangle_{Q}=\dfrac{\left\langle h^{2}\mathbf{1}\left(h\in\left[\frac{1}{\sqrt{M(N)}},1\right]\right)\right\rangle}{P\left(h\in\left[\frac{1}{\sqrt{M(N)}},1\right]\right)}$
We have
[…]
and
[…]
Thus, we have
$\langle h^{2}\rangle_{Q}=\frac{\log M(N)}{M(N)}$
By applying the result to (4) and using $M(N)=(\log N)\cdot K(N)$, we obtain (8).
Appendix B Numerical verification of theoretical results
B.1. Detour distances in ER networks
We verify the derived upper bounds on detour distances in ER networks given by (5) using simulation. Fig. 2 shows the comparison between the simulated detour distances and theoretical bounds in (5) in ER networks for varying number of landmarks $K(N)$. We evaluated the cases where $K(N)=\log N$, $N^{0.5}$ and $N^{0.9}$. In all cases, we observe that the derived upper bound provides very good estimates on the actual detour distances.
Figure 2. Comparison of inter-node distances, landmark detour distances, and upper bound for $K(N)=\log N,N^{0.5},N^{0.9}$ in ER networks.
B.2. Detour distances in BA networks
We verify the derived upper bounds on detour distances in BA networks given by (4). Fig. 3 shows a comparison between the derived upper bounds and the simulated detour distances in BA networks. We observe that the theoretical bound is an excellent match with the simulated distances. In addition, the inter-node distances and the theoretical bounds are quite close to the shortest path distances.
Figure 3. Comparison of inter-node distances, landmark detour distances, and theoretical upper bounds in BA networks.
Appendix C Ablation study
C.1. Graph Clustering Algorithm
We conducted ablation study such that FluidC is replaced by Louvain algorithm (Blondel et al., 2008) which is widely used for graph clustering and community detection. Table 5 shows that our model performs better with FluidC than with Louvain algorithm in all datasets. The proposed hierarchical clustering using FluidC can control the number of clusters to achieve a good trade-off between complexity and performance. However, Louvain algorithm automatically sets the number of clusters, and the number varied significantly over datasets. Moreover, for COLLAB dataset, Louvain algorithm could not be used due to resource issues. Thus, we conclude that FluidC is the better choice in our framework.
Dataset FluidC Louvain
COLLAB 56.24 $\pm$ 0.45 -
DDI 70.97 $\pm$ 6.88 60.79 $\pm$ 8.89
Pubmed 97.56 $\pm$ 0.19 95.83 $\pm$ 0.23
Cora 95.40 $\pm$ 0.21 93.78 $\pm$ 0.15
Citeseer 96.36 $\pm$ 0.17 94.98 $\pm$ 0.12
Facebook 99.69 $\pm$ 0.00 99.30 $\pm$ 0.00
Table 5. Comparison of graph clustering algorithms.
C.2. Node Centrality
For each cluster, the most “central” node should be selected as the landmark. We have used Degree centrality in this paper; however, there are other types of centrality such as Betweenness and Closeness. Betweenness centrality is a measure of how often a given node is included in the shortest paths between node pairs. Closeness centrality is the reciprocal of the sum-length of shortest paths to the other nodes.
Table 6 shows the experimental results comparing Degree, Betweenness and Closeness centralities. The results show that the performances are similar among the centralities. Thus, all the centralities are effective measures for identifying “important” nodes. Degree centrality, however, was better than the other choices.
More importantly, Betweenness and Closeness centralities require full information on inter-node distances, which incurs high computational overhead. In Table 6, we excluded datasets CITATION2 and COLLAB which are too large graphs to compute Betweenness and Closeness centralities. Scalability is crucial for link prediction methods. Thus we conclude that, from the perspective of scalability and performance, Degree centrality is the best choice.
Centrality DDI PubMed Cora Citeseer Facebook
Degree 70.03 $\pm$ 7.02 97.38 $\pm$ 0.34 94.95 $\pm$ 0.18 96.15 $\pm$ 0.19 99.69 $\pm$ 0.00
Betweeness $69.71\pm 6.87$ $97.16\pm 0.19$ $94.69\pm 0.12$ $95.84\pm 0.08$ $99.52\pm 0.00$
Closeness $69.56\pm 5.65$ $96.92\pm 0.20$ $94.43\pm 0.10$ $95.65\pm 0.12$ $99.45\pm 0.00$
Table 6. Link prediction results from landmark selection with different centrality.
Appendix D Actual training time and GPU memory usage
Table 8 and 8 show the comparison and actual training time and GPU memory usage between vanilla GCN with HPLC. Note that HPLC already contains vanilla GCN as a component. Thus, one should attend to the additional resource usage incurred by HPLC in Table 8 and 8. Experiments are conducted with NVIDIA A100 with 40GB memory. We observe that, the additional space and time complexity incurred by HPLC relative to vanilla GCN is quite reasonable, demonstrating the scalability of our framework.
Dataset vanilla GCN HPLC
CITATION2 282856 (ms) 352670 (ms)
COLLAB 1719 (ms) 2736 (ms)
DDI 3246 (ms) 3312 (ms)
PubMed 8859 (ms) 8961 (ms)
Cora 982 (ms) 1022 (ms)
Facebook 1107 (ms) 1243 (ms)
Table 7. Comparison of training time for 1 epoch between vanilla GCN and HPLC.
Dataset vanilla GCN HPLC
CITATION2 29562 (MB) 36344 (MB)
COLLAB 4988 (MB) 7788 (MB)
DDI 2512 (MB) 2630 (MB)
PubMed 2166 (MB) 3036 (MB)
Cora 1458 (MB) 6160 (MB)
Facebook 2274 (MB) 2796 (MB)
Table 8. Comparison of GPU memory usage during training between vanilla GCN and HPLC.
Appendix E Dataset
Dataset statistics is shown in Table 9.
Dataset # Nodes # Edges #$\frac{\#\textrm{Edges}}{\#\textrm{Nodes}}$ Avg. node deg Density Split ratio Metric Cora 2,708 7,986 2.95 5.9 0.0021% 70/10/20 AUC Citeseer 3,327 7,879 2.36 4.7 0.0014% 70/10/20 AUC PubMed 19,717 64,041 3.25 6.5 0.00033% 70/10/20 AUC Facebook 4,039 88,234 21.85 43.7 0.0108% 70/10/20 AUC DDI 4,267 1,334,889 312.84 500.5 14.67% 80/10/10 Hits@20 CITATION2 2,927,963 30,561,187 10.81 20.7 0.00036% 98/1/1 MRR COLLAB 235,868 1,285,465 5.41 8.2 0.0046% 92/4/4 Hits@50
Table 9. Dataset statistics.
Appendix F Hyperparameters
F.1. Number of Clusters
Table 10 shows the performance with varying number of clusters. Specifically, $K=\eta\log N$ where we vary hyperparameter $\eta$. The results show that performance can be improved by adjusting the number of clusters.
$\mathbf{\eta}$ COLLAB DDI PubMed Cora Citeseer Facebook
$1$ 55.52 $\pm$ 0.51 68.27 $\pm$ 6.47 96.81 $\pm$ 0.13 94.34 $\pm$ 0.16 94.70 $\pm$ 0.31 99.53 $\pm$ 0.00
$3$ 55.44 $\pm$ 0.80 68.95 $\pm$ 7.30 96.42 $\pm$ 0.17 94.51 $\pm$ 0.18 96.43 $\pm$ 0.12 99.69 $\pm$ 0.00
$5$ 56.24 $\pm$ 0.45 70.97 $\pm$ 6.88 97.38 $\pm$ 0.34 94.91 $\pm$ 0.11 94.85 $\pm$ 0.14 99.61 $\pm$ 0.00
$7$ 96.89 $\pm$ 0.19 94.95 $\pm$ 0.18 96.15 $\pm$ 0.19 99.63 $\pm$ 0.00
$9$ 97.14 $\pm$ 0.17 94.34 $\pm$ 0.12 95.14 $\pm$ 0.18 99.60 $\pm$ 0.00
$11$ 97.21 $\pm$ 0.24 94.31 $\pm$ 0.19 95.87 $\pm$ 0.13 99.64 $\pm$ 0.00
$15$ 96.76 $\pm$ 0.18 94.40 $\pm$ 0.21 95.36 $\pm$ 0.17 99.65 $\pm$ 0.00
Table 10. Performance of hierarchical graph clustering adjusting hyperparameter $\eta$.
F.2. Miscellaneous
Miscellaneous hyperparameters are shown in Table 11.
Hyperparameter Value Encoder of all plug-in methods GCN Learning rate 0.001, 0.0005 Hidden dimension 256 Number of GNN layers 2, 3 Number of Decoder layers 2, 3 Negative sampling Uniformly Random sampling Dropout 0.2, 0.5 Negative sample rate 1 Activation function ReLU (GNNs), LeakyReLU ($f_{i(k)}$) Loss function BCE Loss Use edge weights False (only binary edge weights) The number of landmarks $O(\log N)$ Optimizer Adam (Kingma and Ba, 2015)
Table 11. Detailed hyperparameters.